OpenStudy (anonymous):

Solve: x + y = –3 x – y = 1 Select one: a. (–1, –2) b. (–3, –4) c. (0, –3) d. (1, 2)

OpenStudy (anonymous):

@eric_d @FriedRice

OpenStudy (anonymous):

it the same as how we both taught you

OpenStudy (anonymous):

Just add both the equations..

OpenStudy (anonymous):

just use elimination method to find your answer

OpenStudy (anonymous):

@FriedRice you carry on... :)

OpenStudy (anonymous):

no you can just do it im helping someone right now

OpenStudy (anonymous):

@eric_d

OpenStudy (anonymous):

ok well if you want our help you must try and we will know if your not.... so if your ready lets start

OpenStudy (anonymous):

what will the first step in the elimination method process

OpenStudy (anonymous):

of course im ready i posted the question

OpenStudy (anonymous):

as how i and eric showed you

OpenStudy (anonymous):

can you show me again please

OpenStudy (anonymous):

well you have the equation x+ y = -3 x - y = 1 right?

OpenStudy (anonymous):

what do you think is the easiest and the fastest way you can get your answer? A.) multiply the first equation by -1 B.)just leave both the equations by itself and just eliminate the y

OpenStudy (anonymous):

actually i rgink the answer us C am i correct

OpenStudy (anonymous):

not even close your just guessing

OpenStudy (anonymous):

shut up

OpenStudy (anonymous):

looooooooser

OpenStudy (anonymous):

hmmmm..... that is pretty cruel.

OpenStudy (anonymous):

just kidding :))))))))

OpenStudy (anonymous):

were friends bro

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well ill just show you all the way for the sake of time. x + y = -3 x - y = 1 2x = -2 x = -1 -1 + y = -3 y = -2 so the answer will be A__________________________A